The area

of an equilateral triangle is 9 √3

square feet. Find a side of the triangle.



1
Expert's answer
2021-02-02T05:00:14-0500

Let "a=" the side of the equilateral triangle. Then


"A=\dfrac{a^2\sqrt{3}}{4}"

Given "Area=9\sqrt{3}\ ft^2."


"\dfrac{a^2\sqrt{3}}{4}=9\sqrt{3}, a>0"

"a^2=36"

"a=6"

The side of the given equilateral triangle is 6 foot.